Solve the following initial value problem

d^2 y/dx^2 + dy/dx -2y = -6 sin2x - 18 cos2x
y(0)=2 , y'(0)=2
1
Expert's answer
2020-04-14T17:39:59-0400

The given differential equation is


"\frac{d^2y}{dx^2}+\frac{dy}{dx}-2y=-6sin(2x)-18cos(2x)"


The auxiliary equation is "m^2+m-2=0."

"\therefore \ m^2+m-2=0"

"\implies m^2+2m-m-2=0"

"\implies m(m+2)-1(m+2)=0"

"\implies (m+2)(m-1)=0"

"\implies m=1,-2" .

"\therefore \ C.F.=Ae^x+Be^{-2x}"

Where ,"C.F.=" Complementary Function.


Now ,

"P.I.=\frac{1}{D^2+D-2}(-6sin2x-18cos2x)"

"=\frac{1}{D^2+D-2}(-6sin2x)+\frac{1}{D^2+D-2}(-18cos2x)""=-6×\frac{1}{-2^2+D-2}sin2x-18×\frac{1}{-2^2+D-2}cos2x""=-6×\frac{1}{D-6}sin2x-18×\frac{1}{D-6}cos2x""=-6×\frac{D+6}{D^2-36}sin2x-18×\frac{D+6}{D^2-36}cos2x"

Multiplying numerator and denomination by "D+6" .


"=-6×\frac{D+6}{-4-36}sin2x-18×\frac{D+6}{-4-36}cos2x"

"=\frac{3}{20}×(D+6)sin2x+\frac{9}{20}×(D+6)cos2x"

"=\frac{3}{20}×(2cos2x+6sin2x)+\frac{9}{20}×(-2sin2x+6cos2x)"

"=cos2x(\frac{3}{10}+\frac{27}{10})+sin2x(\frac{9}{10}-\frac{9}{10})"

"=3cos2x"


Therefore the complete solution is

"y=C.F.+P.I"

"=Ae^x+Be^{-2x}+3cos2x..............(1)"

Again differentiate equation (1) ,we get

"y'(x)=Ae^x-2Be^{-2x}-6sin2x \ ....................(2)"

As given that ,

"y(0)=2 \ and \ y'(0)=2"

So from equation (1) and (2) ,we get

"A+B+3=2"

"\implies A+B=-1 \ ..............(3)"

and "A-2B=2....................(4)"

Subtracting (4) from (3) ,we get

"A+B-(A-2B)=-1-2"

"\implies 3B=-3 \ \implies B=-1"

Again from (3) ,we get

"A-1=-1 \ \implies A=0" .

Therefore equation (1) become

"y=-e^{-2x}+3cos2x" .